elimination
x+3y=14
2x-3y=-8

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)
Answer by ewatrrr(24785)


Hi
solving by elimination (certainly Sets Up for that)
x+3y=14 |green line
2x-3y=-8 |blue line
3x = 6
x = 2
x+3y = 14
2 + 3y = 14
3y = 12
y = 4